OpenStudy (anonymous):

what is the solution to the proportion 4/12=10/x A. 3 B. 4.8 C. 3.3 D. 30

OpenStudy (bohotness):

what do you think the answer is?

OpenStudy (clongoria2):

4 / 12 = 0.33 = 10 / ? = 0.33 Plugin the answers and fine the correct one

OpenStudy (clongoria2):

find*

OpenStudy (anonymous):

3??

OpenStudy (anonymous):

oh wait 3.3

OpenStudy (anonymous):

D. 30 x=10*12/4

OpenStudy (clongoria2):

10 / 3 = 3.33 This is not the correct answer

OpenStudy (clongoria2):

@Srik246 You aren't suppose to give the direct answer

OpenStudy (anonymous):

umm... so is it 30

OpenStudy (clongoria2):

Yes 10/30 = 0.33 which is = to 4/12

OpenStudy (anonymous):

thank you